Good morning distinguished Tutors,i kindly need help with the following question-------"A mining company has 10 000 meters of fencing available .It wants to use the fencing to enclose a rectangular field .One side of the field is bordered by a river .If no fencing is placed on the side next to the river ,what is the largest area that can be enclosed?"
Answer by stanbon(75887)

"A mining company has 10 000 meters of fencing available .It wants to use the fencing to enclose a rectangular field .One side of the field is bordered by a river .If no fencing is placed on the side next to the river ,what is the largest area that can be enclosed?"
------
Sketch the picture
Let length of the rectangle = L
There are two widths; each is W
------
L + 2W = 10,000 meters
----
Area = L*W
-------
Substitute to get:
A = (10,000-2W)(W) = 10,000W - 2W^2
----
Form the derivative and solve for zero
A' = 10,000-4W
10,000-4W = 0
4W = 10,000
W = 2500 (width to get maximum Area)
L = 10,000- 2(2500) = 5000 (length to get maximum Area)
------------------
Cheers,
Stan H.
-----------------